OpenStudy (anonymous):

Whats the solution for this equation? 5w+9=19

OpenStudy (anonymous):

\[5w+9=19\] \[5w+9-9=19-9\] \[5w=10\] \[\frac{10}{5}=w\] \[w=2\]